# Mitigating the Impact of Attribute Editing on Face Recognition

### 12 Mar 2024 v1

@@ Abstract
Facial attribute editing using generative models can impair automated face recognition. This degradation persists even with recent identity-preserving models such as InstantID. To mitigate this issue, we propose two techniques that perform local and global attribute editing. Local editing operates on the finer details via a regularization-free method based on ControlNet conditioned on depth maps and auxiliary semantic segmentation masks. Global editing operates on coarser details via a regularization-based method guided by custom loss and regularization set. In this work, we empirically ablate twenty-six facial semantic, demographic and expression-based attributes altered using state-of-the-art generative models and evaluate them using ArcFace and AdaFace matchers on CelebA, CelebAMaskHQ and LFW datasets. Finally, we use LLaVA, a vision-language framework for attribute prediction to validate our editing techniques. Our methods outperform SoTA (BLIP, InstantID) at facial editing while retaining identity.
@@ 1 Introduction
Facial attribute editing refers to the process of digitally altering semantic cues to change the visual appearance in a face image. Examples may include changing hair color, adding accessories such as eyeglasses, and altering facial expressions; all these go considerably beyond digital re-touching [3, 5] via smartphone applications such as YouCam Makeup and Facetune Editor. Tools for attribute modification are readily available; many are based on conditional generative models (e.g., Cafe-GAN and STGAN, GAN-Edit, Mask-aware editing, Age-GAN and Talk-to-Edit) [20, 24, 25, 35, 40, 17].
@@
Motivation: Although attribute editing models are typically guided by aesthetic or creative pursuits, their impact on face recognition (FR) is relatively under-explored [1]. Automated face matchers demonstrate bias towards semantic attributes and accessories as shown in [37]. Additionally, if digital manipulations are induced, that may adversely impact face recognition. Can generative models be therefore used to deliberately perturb the biometric quality of an image [41]? This question is paramount from the security perspective if digital attribute editing can potentially be used as a passive software-based presentation attack with the intent of evasion (conceal one’s true identity) [18, 28]. The recent, massive proliferation of open-source generative diffusion models that can be used for face synthesis and editing adds urgency to this study [8].
@@
Objective: Concretely, we investigate the impact of a diverse set of semantic, demographic and expression-based attribute editing operations using generative models on automated FR systems. Most generative models rely on MSE, FID or SNR [6] to preserve perceptual quality and not biometric utility. Existing FR systems are equipped to handle low resolution, pose and demographic variations [11, 19] but may falter in presence of digitally induced perturbations. So, we propose techniques that can mitigate the impact of digital attribute editing on FR.
@@
Our Approach: Facial attribute editing is a complex task that can require either minor adaptations locally (e.g., changing hair color) or major alterations globally (e.g., changing age). In both cases, attribute editing should not compromise biometric fidelity. Therefore, we formulate the problem of attribute editing at two levels: (1) local attribute editing, that changes fine facial details. To achieve this, we use segmentation masks to perform the desired edit and depth maps to preserve the extraneous details excluded from the mask. We only need a single facial image to perform local editing. (2) global attribute editing, that changes the coarse facial details. To achieve this, we use a regularization method supervised by identity-preserving loss and utilizes few exemplars of the desired attribute (30 images) and the target individual (10 images). We explore a suite of generative models operating in text-to-image (txt2img) and image-to-image (img2img) modes to conduct our investigation. Tab. 1 outlines the different methods used in this work. See Fig. 0(a) for the overview of the proposed frameworks.
@@ 2 Related Work
Digital retouching [3] and facial attribute editing using generative adversarial networks (GANs) [13, 24, 20, 25, 40, 5, 36, 35, 21, 17, FADE] are highly common. Attribute editing is typically used for creative pursuits and the edited images are evaluated in terms of perceptual quality metrics such as FID and IS. In [1], the authors showed that GAN-based facial attribute editing can significantly degrade automated face recognition performance although the images may be visually perceptible. Image personalization has significantly evolved after the introduction of diffusion models that require a simple text prompt and some images of the target individual (in some cases, even a single image will suffice). The highly non-linear operations performed by the diffusion models may degrade the biometric utility of the generated images, specifically, after semantic editing. Bias in face images generated by diffusion models focus mostly on demographic attributes (age, sex and ethnicity) [2, 29]. Some work propose prompt engineering for de-biasing [7]. ChatFace [42] performs text-driven facial image editing in the semantic latent space of diffusion model This system leverages LLMs to perform complex multi-attribute manipulations through dialogue, similar to Talk-to-Edit [17] which uses interactive facial editing via GANs.
@@ 3 Preliminaries
Table 1: Taxonomy of methods explored in this work.
@@ 6 Results
(a) Prompt: photo of a person with blond hair
(b) Prompt(s): photo of a person with * expression.
@@
Limitations. We observe that the regularization-based method works better on global attributes such as age and facial expression editing while the regularization-free method works better on fine-grained local attribute editing. For expression editing using the CN-IP model, we provide the face mask as conditional input which results in the inpainting model to hallucinate facial details producing random faces. See Fig. 8 for examples of failure cases.
@@ Appendix C Additional examples from CelebA
We present results of images generated using DB-base and DB-prop. on CelebA dataset in Figs. 9 and 10, respectively. We selected SDV2.0 model for DB-base and SDV1.5 model for DB-prop. in this example.

### 09 Apr 2024 v2

@@ Abstract
Through a large-scale study over diverse face images, we show that facial attribute editing using modern generative AI models can severely degrade automated face recognition systems. This degradation persists even with identity-preserving generative models. To mitigate this issue, we propose two novel techniques for local and global attribute editing. We empirically ablate twenty-six facial semantic, demographic and expression-based attributes that have been edited using state-of-the-art generative models, and evaluate them using ArcFace and AdaFace matchers on CelebA, CelebAMaskHQ and LFW datasets. Finally, we use LLaVA, an emerging visual question-answering framework for attribute prediction to validate our editing techniques. Our methods outperform the current state-of-the-art at facial editing (BLIP, InstantID) while improving identity retention by a significant extent.
@@ 1 Introduction
Motivation: Facial attribute editing refers to the process of digitally altering semantic cues to change the visual appearance in a face image. Examples may include changing hair color, adding accessories such as eyeglasses, and altering facial expressions; all these go considerably beyond digital re-touching [3, 5] via smartphone applications such as YouCam Makeup and Facetune Editor. Generative AI-based tools for attribute modification are readily available; many are based on conditional generative models (examples include Cafe-GAN, STGAN, GAN-Edit, Mask-aware editing, Age-GAN, and Talk-to-Edit) [21, 25, 26, 36, 41, 18].
@@
Although attribute editing models are typically guided by aesthetic or creative pursuits, their impact on face recognition (FR) is relatively under-explored [1]. Automated face matchers demonstrate bias towards semantic attributes and accessories as shown in [38]. Additionally, if digital manipulations are induced, that may adversely impact face recognition. Can generative models be therefore used to deliberately perturb the biometric quality of an image? This question is paramount from the security perspective if digital attribute editing can potentially be used as a passive software-based presentation attack with the intent of evasion (conceal one’s true identity) [19, 29]. The recent, massive proliferation of open-source generative diffusion models capable of face synthesis and editing adds urgency to this question [9].
@@
Objective: Our goal will be to investigate the impact of a diverse set of semantic, demographic and expression-based attribute editing operations using generative models on automated FR systems. Most generative models rely on MSE, FID or SNR [7] to preserve perceptual quality and not biometric utility. Existing FR systems are equipped to handle low resolution, pose and demographic variations [12, 20] but may falter in presence of digitally induced perturbations. So, we propose techniques that can mitigate the impact of digital attribute editing on FR.
@@
Our Approach: Facial attribute editing is a complex task that can require either minor adaptations locally (e.g., changing hair color) or major alterations globally (e.g., changing age). In both cases, attribute editing should not compromise biometric fidelity. Therefore, we formulate the problem of attribute editing at two levels: (1) local attribute editing, that changes fine facial details. To achieve this, we use segmentation masks to perform the desired edit and depth maps to preserve the extraneous details excluded from the mask. We only need a single facial image to perform local editing. (2) global attribute editing, that changes the coarse facial details. To achieve this, we use a regularization method supervised by identity-preserving loss and utilizes few exemplars of the desired attribute (30 images) and the target individual (10 images). We explore a suite of generative models operating in text-to-image (txt2img) and image-to-image (img2img) modes to conduct our investigation. Table 1 outlines our different methods. See Fig. 0(a) for the overview of the proposed framework.
@@ 2 Related Work
Digital retouching [3] and facial attribute editing using generative adversarial networks (GANs) [14, 25, 21, 26, 41, 5, 37, 36, 22, 18, 6] are highly common. Attribute editing is typically used for creative pursuits and the edited images are evaluated in terms of perceptual quality metrics such as FID and IS. In [1], the authors showed that GAN-based facial attribute editing can significantly degrade automated face recognition performance although the images may be visually perceptible. Image personalization has significantly evolved after the introduction of diffusion models that require a simple text prompt and some images of the target individual (in some cases, even a single image will suffice). The highly non-linear operations performed by the diffusion models may degrade the biometric utility of the generated images, specifically, after semantic editing. Bias in face images generated by diffusion models focus mostly on demographic attributes (age, sex and ethnicity) [2, 30]. Some work propose prompt engineering for de-biasing [8]. ChatFace [42] performs text-driven facial image editing in the semantic latent space of diffusion model This system leverages LLMs to perform complex multi-attribute manipulations through dialogue, similar to Talk-to-Edit [18] which uses interactive facial editing via GANs.
@@ 3 Preliminaries
Table 1: Taxonomy of methods explored in this work. The top four methods achieve global editing; the bottom three achieve local editing. The methods shaded in blue perform the best in terms of biometric preservation and successful attribute editing.
@@ 6 Results
Limitations. We observe that the regularization-based method works better on global attributes such as age and facial expression editing while the regularization-free method works better on fine-grained local attribute editing. For expression editing using the CN-IP model, we provide the face mask as conditional input which results in the inpainting model to hallucinate facial details producing random faces.
@@ Appendix C Additional examples from CelebA
We present results of images generated using DB-base and DB-prop. on CelebA dataset in Figs. 8 and 9, respectively. We selected SDV2.0 model for DB-base and SDV1.5 model for DB-prop. in this example.
@@
t-SNE analysis. We analyze the AdaFace embedding for the generated images particularly for the attributes that are affecting reliable face recognition. We use t-SNE [tsne] to reduce the 512-D feature vector to 3-D and present the plots for hair color (black hair, blonde hair and brown hair) for InstantID, BLIPDiffusion and CN-IP (ours) in Fig. 10. We observe that BLIPDiffusion forms well-defined clusters for different hair colors while in the case of Instant-ID and CN-IP, the embeddings are evenly distributed. Identity embeddings should not adhere to hair color variations, the tight clustering of embeddings from BLIPDiffusion indicates the contrary. As a result, we observe worse biometric performance from BLIPDiffusion compared to both InstantID and CN-IP.
(a) BLIPDiffusion
(b) InstantID
(c) CN-IP(ours)
Figure 10: Comparison of 3-D t-SNE plots for AdaFace embedding for BLIPDiffusion, InstantID and CN-IP (local). Identity embeddings are expected to be evenly distributed regardless of the hair color variations. However, we note that the ID embeddings from BLIPDiffusion images present relatively distinct pattern for different hair color explaining the loss in biometric fidelity compared to InstantID and CN-IP.
We further compare local attribute editing with global attribute editing using t-SNE in Fig. 11. We observe that even our global attribute editing method is able to maintain evenly distributed identity embeddings across bald, young and male attributes thus, reflecting adherence to biometric features irrespective of different types of semantic and demographic attribute edits.
(a) Bald
(b) Young
(c) Male
Figure 11: Comparison of 3-D t-SNE plots for AdaFace embedding for BLIPDiffusion, InstantID and DB-proposed (global) on bald, young and male attributes. Identity embeddings are expected to be evenly distributed regardless of the variations in hair style, age and digital alteration of sexual cues. We note that our DB-prop. is able to retain such coarse details with better biometric fidelity compared to BLIPDiffusion. Note InstantID defaults to the original image in several cases resulting in identity preservation but fails to perform attribute editing.
Comparison with GAN. We present additional comparisons with Talk-to-Edit [18], an interactive GAN-based facial editing technique. Talk-to-Edit provides fine-grained facial attribute editing via dialog interaction, similar to our approach. The method uses a language encoder to convert the user’s request into an ‘editing encoding’ that encapsulates information about the degree and direction of change of the target attribute, and seeks user feedback to iteratively edit the desired attribute. The authors use the semantic field to preserve attribute localization and an identity-retention loss for maintaining identities. Talk-to-Edit works very well on synthetic images but struggles with real images. We observe that Talk-to-Edit can produce several artifacts (fails to add Bangs in case of black hair, degrades face when asked to increase the intensity of Smiling, fails to remove facial hair when asked with No beard, fails to detect presence of Eyeglasses, and defaults to the original for a majority of face images when asked to change age with Young attribute edit operation.
Failure cases of GAN-based facial attribute editing operation such as Talk-to-Edit [18]. In a majority of cases, it either fails to add attribute thus reconstructing the original image or produces unrealistic artifacts. In Eyeglasses, the method fails to detect glasses in the input image. So, it produces artifacts when trying to remove the glasses in the first edit direction. Similarly, in the No beard example, it adds facial hair.
See Fig. 12 for examples of failure cases.
(a) Prompt: photo of a person with blond hair
(b) Prompt(s): photo of a person with * expression.